a boy throws a ball upward with a velocity 20m/s . the wind imparts a horizontal acceleration of 4m/s^2 to the left.the angle theta (to the vertical) at which the ball must be thrown so that the ball returns to the boy hand ??

Grade:12

1 Answers

ADITI SINGH
7 years ago
ok so what i have so far is, for y: y= (1/2 a t^2) + (Vo cosθ t) (Vcosθ)^2 = (Vocosθ)^2 + 2ay for x: x= (1/2 a t^2) + (Vo sinθ t) (Vsinθ)^2 = (Vosinθ)^2 + 2ax so far how i see it is im going to substitute my 2 x equations and my 2 y equations, then ill have 2 equations left to subtitute which are θ and time.